# SGLC: Semantic Graph-Guided Coarse-Fine-Refine Full Loop Closing for LiDAR SLAM

### v1

@@ Abstract
Loop closing is a crucial component in SLAM that helps eliminate accumulated errors through two main steps: loop detection and loop pose correction. The first step determines whether loop closing should be performed, while the second estimates the 6-DoF pose to correct odometry drift. Current methods mostly focus on developing robust descriptors for loop closure detection, often neglecting loop pose estimation. A few methods that do include pose estimation either suffer from low accuracy or incur high computational costs. To tackle this problem, we introduce SGLC, a real-time semantic graph-guided full loop closing method, with robust loop closure detection and 6-DoF pose estimation capabilities. SGLC takes into account the distinct characteristics of foreground and background points. For foreground instances, it builds a semantic graph that not only abstracts point cloud representation for fast descriptor generation and matching but also guides the subsequent loop verification and initial pose estimation. Background points, meanwhile, are exploited to provide more geometric features for scan-wise descriptor construction and stable planar information for further pose refinement. Loop pose estimation employs a coarse-fine-refine registration scheme that considers the alignment of both instance points and background points, offering high efficiency and accuracy. We evaluate the loop closing performance of SGLC through extensive experiments on the KITTI and KITTI-360 datasets, demonstrating its superiority over existing state-of-the-art methods. Additionally, we integrate SGLC into a SLAM system, eliminating accumulated errors and improving overall SLAM performance. The implementation of SGLC will be released at https://github.com/nubot-nudt/SGLC.
@@
Semantic Graph, Loop Closing, LiDAR SLAM
@@ I Introduction
While many existing methods [1, 2, 3, 4, 5, 6] focus on loop closure detection (LCD), few can estimate closed-loop pose in six degrees of freedom (6-DoF). Some methods incorporate pose estimation with 1-DoF [7, 8, 9] or 3-DoF [10], which may not be sufficient for 6-DoF SLAM systems. Exiting full 6-DoF pose estimation methods, however, are either very time-consuming [11] or have relatively lower accuracy [12].
@@
In summary, our contributions are fourfold: (i) We propose a novel semantic graph-guided full loop closing framework, SGLC, that offers robust loop detection and accurate 6-DoF pose estimation. (ii) We design an effective and efficient outlier pruning method to remove incorrect node correspondences. (iii) We proposed a coarse-fine-refine registration scheme that improves the accuracy and efficiency of pose estimation. (iv) We seamlessly incorporate SGLC into a current SLAM framework, reducing the error in odometry estimations. Extensive evaluation results on KITTI [15] and KITTI-360 [16] datasets support these claims.
@@ II Related Work
Handcrafted-based approaches. These methods typically utilize the geometric features within the point cloud to describe it. In the early stages, Magnusson et al. [17] explore the application of the Normal Distributions Transform (NDT) for surface representation to develop histogram descriptors based on surface orientation and smoothness. The descriptors perform well in different environments and provide the inspiration for subsequent NTD-based methods [18, 19]. To quickly extract descriptors, some methods project point clouds into 2D plane [1, 8, 4, 10, 7] for encoding features. Among them, Scan Context (SC) family [4, 7, 10] is widely used due to its high efficiency and good LCD performance. It converts the raw point cloud into a polar coordinate bird’s eye view (BEV) and encodes the maximum height into the image bins. This yields rotation-invariant ring key descriptors for retrieval and a more detailed 2D image matrix for calculating similarity. Based on that, Wang et al. [4] argue that intensity information of point cloud is effective for loop closing, so they encode the intensity property into SC. Recently, Cui et al. [12] propose BoW3D, a bag of words (BoW)-based method leveraging Link3D features [20], which not only has robust LCD capability but also can estimate 6-DoF pose.
@@
DNN-based approaches. Deep Neural Networks (DNN)-based methods commonly leverage deep neural networks to extract local features and then aggregate them into a global descriptor for retrieval. The early DNN-based work PointNetVLAD [2] is a combination of PointNet [21] for extracting features and NetVLAD [22] for generating the final descriptor. After this, a variety of learning-based loop closing methods begin to proliferate. Liu et al. [23] proposed an adaptive feature extraction module and a graph-based neighborhood aggregation module for enhancing PointNetVLAD. In contrast to them, Komorowski et al. [24] build a sparse voxelized point cloud representation and extract features by sparse 3D convolutions. Chen et al. [9] propose OverlapNet, a range image-based LCD method by estimating image overlap. A subsequent enhanced version, OverlapTransformer [5] is proposed with superior performance and higher efficiency. Cattaneo et al. [11] proposed LCDNet, a robust LCD and 6-DoF pose estimation network validated in various environments, but its high computational complexity makes real-time running challenging. Recently, Luo et al. [6] proposed a lightweight BEV-based NetVALD [22] network that exhibits high descriptor generation efficiency and LCD capability. However, its descriptor dimension far exceeds others, leading to substantial storage consumption, particularly for robots only equipped with low-cost microcontrollers.
@@
Some methods also incorporate semantic information to enhance the descriptor distinctiveness. Kong et al. [3] and Zhu et al. [14] attempt to build semantic graphs for loop closing, while Li  et al. [10] utilize semantics to improve existing descriptors. We also believe semantics are helpful for loop closing as they distinguish different objects in the scene beforehand. However, few semantic-based methods work efficiently and effectively for 6-DoF loop closing.
@@ III-A Semantic Graph
where $N$ is the number of nodes.
Based on the semantic graph, we design the distinctive instance node descriptor for each node, which is utilized for the subsequent robust node matching. We first encode the local relationships of nodes, which can be captured by edges connected to themselves. So we categorize and quantify all edges connected to each node, thereby constructing a histogram-based descriptor. Specifically, for a node $\mathbf{v}\in\mathcal{V}$, we categorize all connected edges into different intervals based on their labels and lengths, and count the occurrences within each interval to generate the descriptor $\mathbf{f}_{l}$.
Due to the $\mathbf{f}_{l}$ only capturing the local typologies within the node neighborhood, it does not account for each node’s global properties, such as their global centrality. To enhance the distinctiveness of the descriptor for robust node matching, we employ eigenvalue decomposition on the graph adjacency matrix $\mathbf{A}$, yielding $\mathbf{A}=\mathbf{Q}\mbox{\boldmath$\Lambda$}\mathbf{Q}^{\text{T}}$. The eigenvalues in $\text{diag}(\mbox{\boldmath$\Lambda$})={\lambda_{1},\ldots,\lambda_{n}}$ are arranged in descending order. The $i$-th row of $\mathbf{Q}$ represents the $i$-th instance node embeddings, which capture the node’s global properties in the graph [27]. Therefore, we use the each row of $\mathbf{Q}$ as part of the instance node descriptor. Considering that the dimensions of the matrix $\mathbf{Q}$ generated by each graph are different, and to unify dimensions of the descriptor, we encode each node using the first $k$ columns of $\mathbf{Q}$, i.e., the eigenvectors corresponding to the largest $k$ eigenvalues, thereby generating a row vector $\mathbf{f}_{g}\in\mathbb{R}^{k}$ for each node. Note that the signs of eigenvectors are arbitrary, hence we use their absolute values. Finally, by concatenating $\mathbf{f}_{l}$, $\mathbf{f}_{g}$, we obtain each instance node’s descriptor $\mathbf{f}\in\mathbb{R}^{D}$ in the graph.
@@ III-B LiDAR Scan Descriptor for Retrival
Based on our devised semantic graph, we can effectively detect the loop closure. Theoretically, we could use local node descriptors to determine graph similarity and find loops directly. However, graph matching using nodes for LCD on a scan-scan basis is highly time-consuming, making it unsuitable for real-time SLAM systems. For fast loop candidate retrieving, we design a novel global descriptor for each LiDAR scan, including the foreground descriptor $\mathbf{F}_{f}$ and background descriptor $\mathbf{F}_{b}$.
@@
Besides $\mathbf{F}_{f}$, which captures the topological relationships between foreground instances, we further design a descriptor to exploit the extensive geometric information in the background point clouds. Inspired by Scan Context [7] encoding the maximum height of the raw point cloud into different bins of a polar BEV image, we leverage semantic information to replace the height data, constructing similar descriptors based on the background point cloud to generate rotation-invariant ring key descriptors. It efficiently encodes the appearance characteristics of different background point clouds, offering a concise portrayal at a minimal computational cost.
Finally, we normalize $\mathbf{F}_{f}$ and $\mathbf{F}_{b}$, and concatenate them to form the LiDAR scan descriptor $\mathbf{F}\in\mathbb{R}^{D^{\prime}}$. This comprehensive descriptor incorporates features from both the foreground semantic graph and the background, thus possessing stronger retrieval capabilities. The LiDAR scan descriptor is utilized to retrieve the multiple similar candidate scans in the database. The similarity between descriptors is computed using the Euclidean distance within the feature space. In real applications, we use the FAISS library [28] to establish the descriptors database and perform fast parallel retrieving.
@@ III-C Geometric Verification
Instance Nodes Matching. For the query scan with semantic graph $\mathcal{G}_{q}$ and the target candidate scan with semantic graph $\mathcal{G}_{t}$, we create an affiliation matrix $\mathbf{I}\in\mathbb{R}^{N\times M}$ based on node descriptor similarity, where $N$ and $M$ are the number of nodes in the query and target graphs, respectively. Each element in $\mathbf{I}$ is calculated as:
@@
$\mathbf{I}_{ij}$ denotes the cost of assigning the $i$-th node in $\mathcal{G}_{q}$ to the $j$-th node in $\mathcal{G}_{t}$. A high cost is assigned to reject a match if theirs node labels are inconsistent or bounding box sizes differ significantly.
@@
To reduce the computational complexity, we propose an efficient outlier pruning scheme based on the local structure of nodes before applying RANSAC. We posit that two nodes of a true positive correspondence should exhibit consistent local geometric structures. For a node $\mathbf{v}$, its neighboring nodes within a certain range can be represented as $\{\mathbf{v}_{j}\}_{j=1}^{K}$, where $K$ is the total number of neighbors. $\mathbf{v}$ and any two of its neighbors can form a triangle, denoted as $\Delta_{j}$, and all local triangles of $\mathbf{v}$ is $\{\Delta_{j}\}_{j=1}^{K(K-1)/2}$. In an ideal scenario, corresponding local geometric triangles of matched nodes should be perfectly aligned. However, practical applications are subject to errors from semantic segmentation and clustering. Therefore, we relax this condition to prevent the elimination of correct correspondences. For the matching of local triangles, we consider two triangles to be consistent when their corresponding sides are of equal length. By doing so, we can effectively eliminate the false node correspondences.
Initial Pose Estimation. Based on pruned correspondences, denoted as $\mathcal{M}_{q}^{{}^{\prime}}$ and $\mathcal{M}_{t}^{{}^{\prime}}$, we can efficiently estimate a relative pose between the query scan and candidate scan based on RANSAC and SVD decomposition [31]. Specifically, in each RANSAC iteration, we randomly select three matching pairs of $\mathcal{M}_{q}^{{}^{\prime}}$ and $\mathcal{M}_{t}^{{}^{\prime}}$ to solve the transformation equation based on SVD decomposition as:
@@
Furthermore, we utilize $\mathbf{T}_{\text{coarse}}$ to align the background point clouds of the query scan and candidate scan, and then calculate the cosine similarity of their background descriptors $\mathbf{F}^{q^{\prime}}_{b}$ and $\mathbf{F}^{t^{\prime}}_{b}$. This additional term provides greater reliability compared to using graph similarity alone. If both $S_{\text{graph}}$ and the background similarity exceed their respective thresholds, the query scan and the candidate scan are considered a true loop closure. We then proceed to estimate their precise pose transformation as follows.
@@
$\mathbf{T}_{\text{coarse}}$ is estimated via sparse graph node matching, while $\mathbf{T}_{\text{icp}}$ is determined by dense instance point registration. This sparse-to-dense registration mechanism enhances pose estimation accuracy in a coarse-to-fine manner. Finally, the relative pose is further refined using point-to-plane constraints for background points. This semantically-guided coarse-fine-refine pose estimation process considers the alignment of both instance points and background points. By separately handling instance points and background points, each alignment stage begins with a favorable initial value, ensuring fast matching and convergence, as well as multi-step checked robust pose estimation.
TABLE I: The performance comparison of F1 max scores and Extended Precision on the KITTI and KITTI-360 datasets.
KITTI KITTI-360
Methods 00 02 05 06 07 08 Mean 0002 0009
DNN-based PointNetVLAD 0.779/0.641 0.727/0.691 0.541/0.536 0.852/0.767 0.631/0.591 0.037/0.500 0.595/0.621 0.349/0.515 0.330/0.510
OverlapNet 0.869/0.555 0.827/0.639 0.924/0.796 0.930/0.744 0.818/0.586 0.374/0.500 0.790/0.637 0.308/0.527 0.739/0.605
LCDNet 0.970/0.847 0.966/0.917 0.969/0.938 0.958/0.920 0.916/0.684 0.989/0.908 0.961/0.869 0.998/0.993 0.988/0.734
OT 0.873/0.800 0.810/0.725 0.837/0.772 0.876/0.809 0.625/0.505 0.667/0.510 0.781/0.687 0.796/0.507 0.879/0.528
BEVPlace 0.960/0.849 0.845/0.819 0.885/0.815 0.895/0.815 0.917/0.687 0.967/0.868 0.912/0.809 0.920/0.504 0.938/0.684
Handcrafted Scan Context 0.750/0.609 0.782/0.632 0.895/0.797 0.968/0.924 0.662/0.554 0.607/0.569 0.777/0.681 0.771/0.554 0.851/0.619
LiDAR-Iris 0.668/0.626 0.762/0.666 0.768/0.747 0.913/0.791 0.629/0.651 0.478/0.562 0.703/0.674 0.704/0.552 0.782/0.580
GOSMatch 0.927/0.651 0.720/0.500 0.840/0.612 0.518/0.521 0.939/0.879 0.908/0.571 0.809/0.622 0.694/0.500 0.766/0.575
SSC 0.939/0.826 0.890/0.745 0.941/0.900 0.986/0.973 0.870/0.773 0.881/0.732 0.918/0.825 0.974/0.780 0.970/0.866
BoW3D 0.977/0.981 0.578/0.704 0.965/0.969 0.985/0.985 0.906/0.929 0.900/0.866 0.885/0.906 0.210/0.560 0.682/0.761
Ours 0.999/0.984 0.888/0.899 0.971/0.972 0.995/0.970 0.998/0.994 0.988/0.980 0.973/0.967 0.932/0.934 0.994/0.978
  • [ $F_{1}$ max / $EP$ ], the best scores are highlighted in bold and the second scores are underlined.

### v2

@@ Abstract
Loop closing is a crucial component in SLAM that helps eliminate accumulated errors through two main steps: loop detection and loop pose correction. The first step determines whether loop closing should be performed, while the second estimates the 6-DoF pose to correct odometry drift. Current methods mostly focus on developing robust descriptors for loop closure detection, often neglecting loop pose estimation. A few methods that do include pose estimation either suffer from low accuracy or incur high computational costs. To tackle this problem, we introduce SGLC, a real-time semantic graph-guided full loop closing method, with robust loop closure detection and 6-DoF pose estimation capabilities. SGLC takes into account the distinct characteristics of foreground and background points. For foreground instances, it builds a semantic graph that not only abstracts point cloud representation for fast descriptor generation and matching but also guides the subsequent loop verification and initial pose estimation. Background points, meanwhile, are exploited to provide more geometric features for scan-wise descriptor construction and stable planar information for further pose refinement. Loop pose estimation employs a coarse-fine-refine registration scheme that considers the alignment of both instance points and background points, offering high efficiency and accuracy. Extensive experiments on multiple publicly available datasets demonstrate its superiority over state-of-the-art methods. Additionally, we integrate SGLC into a SLAM system, eliminating accumulated errors and improving overall SLAM performance. The implementation of SGLC will be released at https://github.com/nubot-nudt/SGLC.
@@
SLAM, Localization
@@ I Introduction
While many existing methods [1, 2, 3, 4, 5, 6] focus on loop closure detection (LCD), few can estimate closed-loop pose in six degrees of freedom (6-DoF). Some methods incorporate pose estimation with 1-DoF [7, 8, 9] or 3-DoF [10, 11], which may not be sufficient for 6-DoF SLAM systems. Existing full 6-DoF pose estimation methods, however, are either very time-consuming [12] or have relatively lower accuracy [13].
@@
In summary, our contributions are fourfold: (i) We propose a novel semantic graph-guided full loop closing framework, SGLC, that offers robust loop detection and accurate 6-DoF pose estimation. (ii) We design an effective and efficient outlier pruning method to remove incorrect node correspondences. (iii) We proposed a coarse-fine-refine registration scheme that improves the accuracy and efficiency of pose estimation. (iv) We seamlessly incorporate SGLC into a current SLAM framework, reducing the error in odometry estimations. Extensive evaluation results on KITTI [16], KITTI-360 [17], Ford Campus [18] and Apollo [19] datasets support these claims.
@@ II Related Work
Handcrafted-based approaches. These methods typically utilize the geometric features within the point cloud to describe it. In the early stages, Magnusson et al. [20] explore the application of the Normal Distributions Transform (NDT) for surface representation to develop histogram descriptors based on surface orientation and smoothness. The descriptors perform well in different environments and provide the inspiration for subsequent NDT-based methods [21, 22]. To quickly extract descriptors, some methods project point clouds into 2D plane [8, 4, 10, 7] for encoding features. Among them, Scan Context (SC) family [4, 7, 10] is widely used due to its high efficiency and good LCD performance. It converts the raw point cloud into a polar coordinate bird’s eye view (BEV) and encodes the maximum height into the image bins. This yields rotation-invariant ring key descriptors for retrieval and a more detailed 2D image matrix for calculating similarity. Based on that, Jiang et al. [11] employed a similar projection-based approach, but it generates multi-layered BEV images, which can effectively reduce the loss of height information. Recently, Cui et al. [13] propose BoW3D, leveraging Link3D features [23] to build bag of words and estimate 6-DoF pose.
@@
Deep Neural Networks (DNN)-based approaches. These methods leverage networks to extract local features and then aggregate them into a global descriptor for retrieval. The early DNN-based work PointNetVLAD [2] is a combination of PointNet [24] for extracting features and NetVLAD [25] for generating the final descriptor. After this, a variety of learning-based loop closing methods begin to proliferate. Liu et al. [26] proposed an adaptive feature extraction module and a graph-based neighborhood aggregation module for enhancing PointNetVLAD. In contrast to them, Komorowski et al. [27] build a sparse voxelized point cloud representation and extract features by sparse 3D convolutions. Chen et al. [9] propose OverlapNet, a range image-based LCD method by estimating image overlap. A subsequent enhanced version, OverlapTransformer [5] is proposed with superior performance and higher efficiency. Cattaneo et al. [12] proposed LCDNet, a robust LCD and 6-DoF pose estimation network validated in various environments, but its high computational complexity makes real-time running challenging. Recently, Luo et al. [6] proposed a lightweight BEV-based network showing high efficiency and LCD capability.
@@
Some methods also incorporate semantic information to enhance the descriptor distinctiveness. Kong et al. [3, 28] and Zhu et al. [15] attempt to build semantic graphs for loop closing, while Li  et al. [10] utilize semantics to improve existing descriptors. Building upon LCDNet [12], Arce et al. [29] integrate semantic information into the network only during the training phase, which improves loop closing performance and makes the model versatile. We agree that semantics are beneficial for loop closing, as they help differentiate objects in the scene. However, few semantic-based methods have proven both efficient and effective for 6-DoF loop closing.
@@ III-A Semantic Graph
Based on the semantic graph, we design instance node descriptors for the subsequent robust node matching. We first encode the local relationships of nodes by using connected edges. These edges are categorized and quantified to construct a histogram-based descriptor. Specifically, for a node $\mathbf{v}\in\mathcal{V}$, we categorize all connected edges into different intervals based on their labels and lengths, and count the occurrences within each interval with length $d_{i}$ to generate the descriptor $\mathbf{f}_{l}$.
Since $\mathbf{f}_{l}$ only captures the local typologies within the node neighborhood, it does not contain node global properties, such as their global centrality. To enhance the distinctiveness of the descriptor for robust node matching, we employ eigenvalue decomposition on the graph adjacency matrix $\mathbf{A}$, yielding $\mathbf{A}=\mathbf{Q}\mbox{\boldmath$\Lambda$}\mathbf{Q}^{\text{T}}$. The eigenvalues in $\text{diag}(\mbox{\boldmath$\Lambda$})={\lambda_{1},\ldots,\lambda_{n}}$ are arranged in descending order. The $i$-th row of $\mathbf{Q}$ represents the $i$-th instance node embeddings, which capture the node’s global properties in the graph [32]. Therefore, we use the each row of $\mathbf{Q}$ as part of the instance node descriptor. Considering that the dimensions of the matrix $\mathbf{Q}$ generated by each graph are different, and to unify dimensions of the descriptor, we encode each node using the first $k$ columns of $\mathbf{Q}$, i.e., the eigenvectors corresponding to the largest $k$ eigenvalues, thereby generating a row vector $\mathbf{f}_{g}\in\mathbb{R}^{k}$ for each node. Note that the signs of eigenvectors are arbitrary, hence we use their absolute values. Finally, by concatenating $\mathbf{f}_{l}$, $\mathbf{f}_{g}$, we obtain each instance node’s descriptor $\mathbf{f}\in\mathbb{R}^{D}$ in the graph.
@@ III-B LiDAR Scan Descriptor for Retrival
Based on our semantic graph, our method effectively detects the loop closure. Theoretically, we could use local node descriptors to determine graph similarity and find loops directly. However, graph matching using nodes for LCD on a scan-scan basis is time-consuming, making it unsuitable for real-time SLAM systems. For fast loop candidate retrieving, we design a novel global descriptor for each LiDAR scan, including the foreground descriptor $\mathbf{F}_{f}$ and background descriptor $\mathbf{F}_{b}$.
@@
Besides $\mathbf{F}_{f}$, which captures the topological relationships between foreground instances, we further design a descriptor to exploit the extensive geometric information in the background point clouds. Inspired by Scan Context [7] encoding the maximum height of the raw point cloud into different bins of a polar BEV image, we leverage semantic information to replace the height data, constructing similar descriptors based on the background point cloud to generate rotation-invariant ring key descriptors. This descriptor efficiently encodes the appearance characteristics of different background point clouds, offering a concise portrayal at a minimal computational cost.
Finally, we perform L2-normalization on $\mathbf{F}_{f}$ and $\mathbf{F}_{b}$ to enhance the robustness, and concatenate them to form the LiDAR scan descriptor $\mathbf{F}\in\mathbb{R}^{D^{\prime}}$. This comprehensive descriptor incorporates features from both the foreground semantic graph and the background, thus possessing stronger retrieval capabilities. The LiDAR scan descriptor is utilized to retrieve the multiple similar candidate scans in the database. The similarity between descriptors is computed using the Euclidean distance within the feature space. In real applications, we use the FAISS library [33] to establish the descriptors database and perform fast parallel retrieving.
@@ III-C Geometric Verification
Instance Nodes Matching. For the query scan with semantic graph $\mathcal{G}_{q}$ and the target candidate scan with semantic graph $\mathcal{G}_{t}$, we create an affiliation matrix $\mathbf{I}\in\mathbb{R}^{N\times M}$ based on node descriptor similarity, where $N$ and $M$ are the number of nodes in the query and target graphs, respectively. Each element $\mathbf{I}_{ij}$ is the cost of assigning the $i$-th node in $\mathcal{G}_{q}$ to the $j$-th node in $\mathcal{G}_{t}$ as:
@@
A high cost is assigned to reject a match if theirs node labels are inconsistent or bounding box sizes differ significantly, i.e., the difference in each dimension of the box beyond $d_{b}$.
@@
To reduce the computational complexity, we propose an efficient outlier pruning scheme based on the local structure of nodes before applying RANSAC. We posit that two nodes of a true positive correspondence should exhibit consistent local geometric structures. For a node $\mathbf{v}$, its neighboring nodes within a certain range $d_{l}$ can be represented as $\{\mathbf{v}_{j}\}_{j=1}^{K}$, where $K$ is the total number of neighbors. $\mathbf{v}$ and any two of its neighbors can form a triangle, denoted as $\Delta_{j}$, and all local triangles of $\mathbf{v}$ is $\{\Delta_{j}\}_{j=1}^{K(K-1)/2}$. In an ideal scenario, corresponding local geometric triangles of matched nodes should be perfectly aligned. However, practical applications are subject to errors from semantic segmentation and clustering. Therefore, we relax this condition to prevent the elimination of correct correspondences. For the matching of local triangles, we consider two triangles to be consistent when their corresponding sides are of equal length. Once the number of successfully matched local triangles between two nodes exceeds $t_{m}$, they are considered as true node correspondences.
Initial Pose Estimation. Based on pruned correspondences, denoted as $\mathcal{M}_{q}^{{}^{\prime}}$ and $\mathcal{M}_{t}^{{}^{\prime}}$, we estimate a relative pose between the query and candidate scans based on RANSAC and SVD decomposition [36]. In each RANSAC iteration, we randomly select three matching pairs of $\mathcal{M}_{q}^{{}^{\prime}}$ and $\mathcal{M}_{t}^{{}^{\prime}}$ to solve the transformation equation based on SVD decomposition as:
@@
Furthermore, we use $\mathbf{T}_{\text{coarse}}$ to align the background point clouds of the query and candidate scans, and then calculate the cosine similarity of their background descriptors $\mathbf{F}^{q^{\prime}}_{b}$ and $\mathbf{F}^{t^{\prime}}_{b}$. This provides greater reliability compared to using graph similarity alone. If both $S_{\text{graph}}$ and the background similarity exceed thresholds $t_{g}$ and $t_{b}$, the query scan and the candidate scan are considered a true loop closure. We then proceed to estimate their precise pose transformation as follows.
@@
$\mathbf{T}_{\text{coarse}}$ is estimated via sparse node matching, while $\mathbf{T}_{\text{icp}}$ is determined by dense instance points registration. This sparse-to-dense registration mechanism enhances pose estimation accuracy in a coarse-to-fine manner. Finally, the pose is further refined using point-to-plane constraints for background points. This semantically-guided coarse-fine-refine pose estimation process considers the alignment of both instance points and background points. By separately handling instance points and background points, each alignment stage begins with a favorable initial value, ensuring fast matching and convergence, as well as multi-step checked robust pose estimation.